### v1

@@ 4.3 Memorization Gaps in GFlowNets
We run the setup explained in §3.2 and Table 1; results are shown in Figures 6, with results involving noise corruption [^6] This is another setting considered by Zhang et al. (2017; 2021) to achieve de-structuring of the form $s\!\perp\!\!\!\!\perp\tilde{R}(s)$. (corrupted constant reward) and online trained GFlowNets included in §D.3. For readers not familiar with the results of Zhang et al. (2021)’s work, a typical observation in the memorization setup is for the training loss to initially plateau, until some phase change occurs (one can imagine parameters have self-arranged to create separate linear regions around each input point) and the training loss goes to 0. We observe a similar phenomenon here, regressing to the original $R$ and $P_{F}$ induce curves with no plateau, while regressing to $\tilde{R}$ and $P_{F}^{\mathrm{random}}$ initially have very flat plateaus until a phase change occurs. This shows a very clear memorization gap, which was expected.
The most interesting result here is where $P_{F}$ is only partially de-structured, and we regress to the true $P_{F}$ of a shuffled reward $\tilde{R}$ and retain the environment structure information. We observe that while there is no flat plateau, (a) $P_{F}$ becomes harder to fit and (b) there remains a phase change whereby learning accelerates after some slow period. These results support Hypothesis 2; $P_{F}$ captures structure of both the reward function and the environment, since removing the structure of either induces memorization gaps.
@@
The existence of clear memorization gaps, when removing reward structure but maintaining environment structure, suggests that there is informative structure in flow functions and flow policies.
This finding implies that state flows may play a mechanistic role in why GFlowNets generalize to unseen states. In §D.3 we show an extension of these results where we consider rewards corrupted with noise (rather than shuffled). We also show similar results for sequences and grids (see §E.3).
@@ 4.4 Generalization in Offline and Off-policy Training of GFlowNets
We run the setup described in §3.3 and Table 2; results on the neighbors task are shown in Figure 7(a) and Figure 7(b). We include full results for all tasks in §D.4. First, we observe that GFlowNets trained in this regime are dependent on environment/task difficulty, the choice of ${\mathbb{P}}_{{\mathcal{X}}}$, and on the evaluation metric. For example, the best performing choice for ${\mathbb{P}}_{{\mathcal{X}}}$ on the counting task is typically $x\sim{\mathcal{U}}$ or $x\sim\log R(x)$, while this is not necessarily the case for neighbors and cliques (see §D.4 Figure 14).
Second, we observe that when training $P_{F}(s^{\prime}|s;\theta)$ offline and off-policy in the full data (no test set) regime, convergence rate depends on the choice of ${\mathbb{P}}_{{\mathcal{X}}}$ (see Figure 7(a)), sometimes to a catastrophic degree. We see this for our usual distributional metrics for approximating $p(x)$ [^7] It is possible that the drawbacks of this training regime overwhelm our ability to make meaningful conclusions on the choice of ${\mathbb{P}}_{{\mathcal{X}}}$ for offline training on the neighbors and cliques tasks.. However, when considering the MAE between $\log R(s)$ the approximated $\log\hat{R}(s)$, we do not observe this lack of convergence. We can measure this because GFlowNets learn to implicitly predict the reward: $\log\hat{R}(s)=\log F(s;\theta)+\log P_{F}(s_{f}|s;\theta)$. This suggests that although training $P_{F}(s^{\prime}|s;\theta)$ offline and off-policy may be tricky, learning $F(s;\theta)$ jointly with $P_{F}(s^{\prime}|s;\theta)$ might be beneficial. This is supported by observing a reasonable estimate for $\log\hat{R}(s)$ even when $P_{F}(s^{\prime}|s;\theta)$ fails to adequately model $p(x)$ (the model is learning something).
We investigate this finding deeper by considering the off-policy interpolation setup and find that for the more difficult tasks (neighbors and cliques) there exists an apparent trade-off between on-policy and off-policy training when estimating $p(x)$ versus $R$ (see §D.4 Figure 15). In particular, when sampling on-policy (from $P_{F}(s^{\prime}|s;\theta)$, $\alpha=0$), approximating $p(x)$ is reasonable. In contrast, approximating $R$ is best when sampling uniformly off-policy (from $P_{{\mathcal{U}}}(s^{\prime}|s)$, $\alpha=1$) even though error for $p(x)$ high.
@@
Deviating from the on-policy training distribution induced by $P_{F}(s^{\prime}|s;\theta)$ can affect the ability of GFlowNets’ to learn $p(x)$.
This finding implies that training GFlowNets offline and off-policy may come with challenges depending on the task. However, we observe that the model could be learning some informative structure in $F(s)$ that could help it generalize (seen by reasonable estimates for $\log\hat{R}(s)$). We test this hypothesis in our second offline and off-policy training setup, and we consider performance on an unseen test (see Figure 7(b)). We observe that indeed offline and off-policy trained GFlowNets reasonably approximate $R$ on unseen states even when failing to adequately model $p(x)$.
@@
Learning $F(s)$ is informative to models, who implicitly predict reasonable rewards, even when $P_{F}$ is unreasonable.
These results support Hypothesis 1; training GFlowNets online and on-policy is ideal, where $P_{F}(s^{\prime}|s;\theta)$ converges to sampling trajectories proportional to $R$. We repeat this experiment (with a left out test set) for the sequence and hypergrid environments. Interestingly, we observe that in these environments $p(x)$ is adequately modelled in the offline and off-policy training regime (see §E.4). This could in-part be due to the reduced complexity of these environments, which we establish by comparing the difficulty of our tasks to the difficulty of fitting the constant reward (see §E.1).
@@ 5 Conclusion
In this work, we conducted an empirical investigation into the generalization behaviours of GFlowNets. We introduced a set of graph generation tasks of varying difficulty to benchmark and measure GFlowNet generalization performance. Our findings support existing hypothesized mechanisms of generalization of GFlowNets as well as add to our understanding of why GFlowNets generalize. We found that GFlowNets inherently learn to approximate functions which contain structure favourable for generalization. In addition, we found that the reward implicitly learned by GFlowNets is robust to changes in the training distribution, but that GFlowNets are sensitive to being trained off-policy. We discuss limitations and future directions of this work in §B.
Impact Statement
The outcomes of the research conducted in this work have practical implications for generative modelling in applications that include drug discovery, material design, and general combinatorial optimization in commercial settings. Because of this, we recognize the importance of considering safety and alignment in these closely related domains. We believe that advancements in GFlowNet research could result in models with improved generalization, in turn leading to models that are easier to align. Another important and useful application of GFlowNets is in the causal and reasoning domains. We believe that advancements in these areas may lead to more interpretable, easier to understand, and safer models. Lastly, we highlight that the applications of our findings for the advancements of GFlowNets may rely on building upon the hypotheses put forward in this work.
@@ Acknowledgments
The bulk of this research was done at Valence Labs as part of an internship, using computational resources there. The authors are grateful to Yoshua Bengio, Berton Earnshaw, Dinghuai Zhang, Johnny Xi, Jason Hartford, Philip Fradkin, Cristian Gabellini, Julien Roy, and the Valence Labs team for fruitful discussions and feedback. In addition, we acknowledge funding from the Natural Sciences and Engineering Research Council of Canada and the Vector Institute.
@@ Appendix B Limitations and Future Work
Limitations:
We have shown that GFlowNets tend to generalize when learning unnormalized probability mass functions for approximating $p(x)$ over discrete spaces. In particular, we have investigate the generalization behaviours of GFlowNets in the context of distributional errors for $p(x)$. Because of the combinatorial nature of computing $p(x)$ exactly, and likewise computing $p(x;\theta)$ from $P(s^{\prime}|s;\theta)$, we are limited in the fact that we need to work within reasonably sized discrete spaces (as those we have proposed) to study GFlowNet generalization for approximating $p(x)$. Secondly, we don’t explicitly test generalization in online and on-policy trained GFlowNets. This in part a consequence of requiring an environment and state space large enough such that a sufficient quantity of unseen states can be produced, while also being able to purposely hide the visited states from parameters updates of the GFlowNet. Lastly, although our proposed problems and environments have structure that induces difficult and interesting generalization tasks, they still may remain far from the structure of the real world.
Future Directions:
@@ C.1.1 Reward functions
        total_correct += int(num_diff_colr % 2 == 0) - (1 if num_diff_colr == 0 else 0)
@@
counting simply counts the number of red and green nodes, red nodes being ”worth” more, and again modifies this count in order to produce a nice 0 to -10 log-reward distribution.
@@
Figure 8: Distribution of log rewards for the given tasks: (left) counting, (middle) neighbors, and (right) cliques.
@@ C.2.2 Bit Sequences:
Auto-regressive sequence environments have been used in a variety of past GFlowNet work Malkin et al. (2022); Madan et al. (2022); Pan et al. (2023). In this work, we consider a bit sequence environment akin to that originally introduced by Malkin et al. (2022). We consider a reward of the form $R(x)=\exp(\frac{-\min_{m\in{\mathcal{M}}}d(x,m))}{l}\times 10)$, where $d$ is the Levenshtein edit distance between an input sequence $x$ and the closest “mode” sequence $m$, ${\mathcal{M}}\subset{\mathcal{X}}$ is the set of mode-sequences, and $l$ is the max sequence length. We select $|{\mathcal{M}}|=60$ sequences from ${\mathcal{M}}$ uniformly at random given the set ${\mathcal{X}}$ of all possible bit sequences up to length $l=15$ to be the “mode” sequences (we do this once at the start of each run). This yields a discrete state space of $65,535$ states, comparable in size to our graph generation environment.
@@ C.3 Monotonic reward transformation
We consider a monotonic transform of the form ${\mathcal{H}}_{\gamma}(\log R(x))=\log R(x)e^{(-\gamma\log R(x))}$. As described earlier, the transform ${\mathcal{H}}_{\gamma}$ maps $\log R^{\prime}(s)={\mathcal{H}}_{\gamma}(\log R(s)),\forall s\in{\mathcal{S}}$ such that $\log R^{\prime}(s)$ is skewed towards higher reward states as a function of some parameter $\gamma$ (see §C.3 for further details. Note, ${\mathcal{H}}_{\gamma}$ is monotonic for values of $\log R(x)\leq 0$ and $\gamma\geq 0$. In our graph and sequence generations tasks, $\log R(x)\leq 0,\forall x\in{\mathcal{X}}$, hence we are able to use ${\mathcal{H}}_{\gamma}$ as a monotonic transform for $\log R(x)$. We show an example of ${\mathcal{H}}_{\gamma}$ applied to the cliques task in Figure 9 for various value sof $\gamma$.
@@ C.4 Evaluation metrics
To measure the generative modelling performance of GFlowNets we consider the Jensen-Shannon (JS) divergence and the MAE between the learned $\log p(x;\theta)$ and the true $\log p(x)$. For JS divergence, we compute:
@@
where $Q=\frac{1}{2}(p(x)+p(x;\theta))$. This reduces to
$\text{JS}(p(x),p(x;\theta))=\frac{1}{2}\sum_{x\in{\mathcal{X}}}\left(p(x;\theta)(\log p(x;\theta)-\log p(x))+p(x)(\log p(x)-\log p(x;\theta))\right).$ (4)
For the MAE, we simply take the absolute error $|\log p(x;\theta)-\log p(x)|$ and average over the cardinality of the state space.
@@ C.5 Model architectures
For all graph experiments we use a modified graph transformer (Veličković et al., 2017; Shi et al., 2020). On top of the normal attention mechanism, we augment the input of each layer with the output of one round of message passing (using layers from the work of Li et al., 2020) with a sum aggregation–we found that this was useful in tasks where counting was required. We use 8 layers with 128-dimensional embeddings and 4 attention heads; we use this architecture after having tried different numbers of layers and embeddings in order to validate our task; this is coincidentally represented in Fig. 1.

### v2

@@ 4.3 Memorization Gaps in GFlowNets
Figure 4: Memorization gap training curves for counting, neighbors, and cliques tasks. Maintaining flow structure in the learning problem (learning $P_{F}$ under shuffled $\tilde{R}$) induces a smaller memorization gap relative to the fully de-structured setting. See Table 3 for reference of experimental setup.
We run the setup explained in §3.2 and Table 3; results are shown in Figure 4. Additional results with $R(s)\sim\mathcal{N}(\mu,\sigma)$ and online trained GFlowNets are in §D.5. A typical observation in the memorization gap setup is for the training loss to initially plateau, until some phase change occurs (one can imagine parameters have self-arranged to create separate linear regions around each input point) and the training loss goes to 0. We se similar trends here, regressing to $R$ and $P_{F}$ induce curves with no plateau, while regressing to $\tilde{R}$ and $P_{F}^{\mathrm{random}}$ initially have very flat plateaus until a phase change occurs. This shows a very clear memorization gap, which was expected.
Most interesting is where $P_{F}$ is partially de-structured, and we regress to the true $P_{F}$ of a shuffled reward $\tilde{R}$ and retain the environment structure. While there is no plateau, $P_{F}$ becomes harder to fit, and there remains a phase change. These results support Hypothesis 2; $P_{F}$ captures structure of both the reward function and the environment, since removing either induces memorization gaps.
@@
The existence of clear memorization gaps, when removing reward structure but maintaining environment structure, upholds that environment structure facilitates generalization when learning flow functions and flow policies.
This finding implies that state flows may play a mechanistic role in why GFlowNets generalize to unseen states. We also show similar results for sequences and grids (see §E.3).
@@ 4.4 Generalization in Offline and Off-policy Training of GFlowNets
We run the offline training setup described in §3.3 and Table 4; results on the neighbors task are shown in Figures 5(a) and 5(b). Results for all tasks are in §D.6. First, we observe that performance is dependent on environment/task difficulty, the choice of ${\mathbb{P}}_{{\mathcal{X}}}$, and on the evaluation metric. For example, the best choice for ${\mathbb{P}}_{{\mathcal{X}}}$ on the counting task is typically $x\sim{\mathcal{U}}$ or $x\sim\log R(x)$, while this is not always the case for neighbors and cliques (see §D.6 Figure 13).
Second, we observe that when training $P_{F}(s^{\prime}|s;\theta)$ offline and off-policy in the full data (no test set) regime, convergence rate heavily depends on the choice of ${\mathbb{P}}_{{\mathcal{X}}}$ (see Figure 5(a)), sometimes to a catastrophic degree. We see this for our usual distributional metrics for approximating $p(x)$ [^7] It is possible that the drawbacks of this training regime overwhelm our ability to make meaningful conclusions on the choice of ${\mathbb{P}}_{{\mathcal{X}}}$ for offline training on the neighbors and cliques tasks.. However, when considering the MAE between $\log R(s)$ the approximated $\log\hat{R}(s)$, we do not observe this lack of convergence (GFlowNets implicitly learn to predict the reward: $\log\hat{R}(s)=\log F(s;\theta)+\log P_{F}(s_{f}|s;\theta)$). Although training $P_{F}(s^{\prime}|s;\theta)$ offline or off-policy may be tricky, learning $F(s;\theta)$ jointly with $P_{F}(s^{\prime}|s;\theta)$ might be beneficial: we observing a reasonable estimate for $\log\hat{R}(s)$ even when $P_{F}(s^{\prime}|s;\theta)$ fails to adequately model $p(x)$—the model is learning something.
We then run the off-policy interpolations (as described in §3.3), with $P_{\alpha}=(1-\alpha)P_{F}+\alpha P_{{\mathcal{U}}}$ and plot the results in Figure 6. We observe a trade-off between on-policy and off-policy training when estimating $p(x)$ versus $R$. We find that for harder tasks (neighbors, cliques), making the data more off-policy ($\alpha=1$) makes $p(x;\theta)$ become worse and $\hat{R}$ better. When sampling on-policy (from $P_{F}(s^{\prime}|s;\theta)$, $\alpha=0$), approximating $p(x)$ is reasonable, but approximating $R$ is best when sampling uniformly off-policy (from $P_{{\mathcal{U}}}(s^{\prime}|s)$, $\alpha=1$).
@@
Deviating from the on-policy training distribution induced by $P_{F}(s^{\prime}|s;\theta)$ can negatively affect the ability of GFlowNets’ to learn $p(x)$.
This implies that training GFlowNets offline or off-policy may come with challenges. However, the model could be learning some informative structure in $F(s)$ that helps it generalize (seen by a reasonable $\log\hat{R}(s)$). We test this hypothesis in our second offline and off-policy training setup, and consider performance on a test set (Figure 5(b)). We observe that offline and off-policy trained GFlowNets reasonably approximate $R$ on unseen states even when failing to adequately model $p(x)$.
@@
Learning $F(s)$ can facilitate models learning to implicitly predict the reward on unseen states, even when $P_{F}$ does not adequately approximate $p(x)$ on these states.
These results support Hypothesis 1; training GFlowNets online and on-policy is ideal, where $P_{F}(s^{\prime}|s;\theta)$ converges to sampling in proportion to $R$. We repeat these experiments for the sequence and hypergrid environments, where we observe that $p(x)$ is adequately modelled in the offline and off-policy cases (see §E.4). This may be due to the lesser complexity of these environments, which we establish by comparing the difficulty of our tasks to that of fitting a constant reward (see §E.1).
@@ 5 Conclusion
In this work, we conducted an empirical investigation into the generalization behaviours of GFlowNets. We introduced a set of graph generation tasks of varying difficulty to benchmark and measure GFlowNet generalization performance. Our findings support existing hypothesized mechanisms of generalization of GFlowNets as well as add to our understanding of why GFlowNets generalize. We found that GFlowNets inherently learn to approximate functions which contain structure favourable for generalization. In addition, we found that the reward implicitly learned by GFlowNets is robust to changes in the training distribution, but that GFlowNets are sensitive to being trained off-policy.
Discussion:
Our objective in this work was to take a scientific approach of starting from existing set of hypotheses and testing them through a more comprehensive set of benchmark tasks. For example, many existing works claim that GFlowNets can effectively operate in the off-policy regime. However, through our investigation, we show that GFlowNets are not as robust as the tone of those claims would suggest. We demonstrated this by introducing a novel graph-based benchmark environment where task difficulty can be easily varied while also varying the off-policy sampling distribution. We repeated our empirical investigation using standard environments which are used in GFlowNet literature (i.e. the hypergrid and sequence environments). Interestingly at times, and particularly for the off-policy results, there is a discrepancy of the observed behavior of GFlowNet generalization performance between the hypergrid and sequence environments with performance on our graph-based environment. This result suggests the importance of taking into consideration the choice of benchmark environment when developing novel GFlowNet methods.
Our intention was to systematically test existing hypotheses regarding GFlowNets and their ability to generalize to unseen area of the state space. In testing these hypotheses, we confirmed existing knowledge while also finding contradictory outcomes from what has been previously understood. For example, while Observations 4.1-4.3 confirmed our existing received knowledge regarding GFlowNets and their generalization capabilities, Observations 4.4 and 4.5 were unexpected results. We believe these are interesting findings that warrant further investigation and could for example lead to improvements in off-policy training and exploration algorithms for GFlowNets.
Limitations:
We have shown that GFlowNets tend to generalize when learning unnormalized probability mass functions for approximating $p(x)$ over discrete spaces. In particular, we investigated the generalization behaviours of GFlowNets in the context of distributional errors for $p(x)$. Because of the combinatorial nature of computing $p(x)$ exactly, and likewise computing $p(x;\theta)$ from $P(s^{\prime}|s;\theta)$, we are limited in the fact that we need to work within reasonably sized discrete spaces (as those we have proposed) to study GFlowNet generalization for approximating $p(x)$. Secondly, we don’t explicitly test generalization in online and on-policy trained GFlowNets. This in part a consequence of requiring an environment and state space large enough such that a sufficient quantity of unseen states can be produced, while also being able to purposely hide the visited states from parameters updates of the GFlowNet. Lastly, although our proposed problems and environments have structure that induces difficult and interesting generalization tasks, they still may remain far from the structure of the real world.
Future Directions:
@@ Acknowledgments
The bulk of this research was done at Valence Labs as part of an internship, using the computational resources there. The authors are grateful to Yoshua Bengio, Berton Earnshaw, Dinghuai Zhang, Johnny Xi, Jason Hartford, Philip Fradkin, Cristian Gabellini, Julien Roy, and the Valence Labs team for fruitful discussions and feedback. In addition, we acknowledge funding from the Natural Sciences and Engineering Research Council of Canada and the Vector Institute.
@@ Appendix A Vocabulary
  • True flows, we occasionally refer to true flows. What we mean by that is the exact calculation of $F(s)$ or $F(s\to s^{\prime})$ as a function of the reward $R(s)$ and of $P_{B}$. Note that for a fixed $P_{B}$ there is a unique solution to $F$ (Bengio et al., 2021). When the reward is corrupted, we talk about the true flow of the corrupted reward $\tilde{R}$, i.e. $F(s)$ is the exact calculation of the flow through $s$ but for terminal (sink) flows set to $\tilde{R}$.
@@ Appendix B Limitations and Future Work
Impact Statement
The outcomes of the research conducted in this work have practical implications for generative modeling in applications that include drug discovery, material design, and general combinatorial optimization in commercial settings. Because of this, we recognize the importance of considering safety and alignment in these closely related domains. We believe that advancements in GFlowNet research could result in models with improved generalization, in turn leading to models that are easier to align. Another important and useful application of GFlowNets is in the causal and reasoning domains. We believe that advancements in these areas may lead to more interpretable, easier to understand, and safer models. Lastly, we highlight that the applications of our findings for the advancements of GFlowNets may rely on building upon the hypotheses put forward in this work.
@@ C.1.1 Reward functions
        total_correct += int(num_diff_colr %
@@
counting simply counts the number of red and green nodes, red nodes being "worth" more, and again modifies this count in order to produce a nice 0 to -10 log-reward distribution.
@@
Figure 7: Distribution of normalized count of log-rewards over the state space for the given tasks: (left) counting, (middle) neighbors, and (right) cliques.
@@ C.2.2 Bit Sequences:
Auto-regressive sequence environments have been used in a variety of past GFlowNet work (Malkin et al., 2022; Madan et al., 2022; Pan et al., 2023). In this work, we consider a bit sequence environment akin to that originally introduced by Malkin et al. (2022). We consider a reward of the form $R(x)=\exp(\frac{-\min_{m\in{\mathcal{M}}}d(x,m))}{l}\times 10)$, where $d$ is the Levenshtein edit distance between an input sequence $x$ and the closest “mode” sequence $m$, ${\mathcal{M}}\subset{\mathcal{X}}$ is the set of mode-sequences, and $l$ is the max sequence length. We select $|{\mathcal{M}}|=60$ sequences from ${\mathcal{M}}$ uniformly at random given the set ${\mathcal{X}}$ of all possible bit sequences up to length $l=15$ to be the “mode” sequences (we do this once at the start of each run). This yields a discrete state space of $65,535$ states, comparable in size to our graph generation environment.
@@ C.3 Monotonic reward transformation
We consider a monotonic transform of the form $\mathcal{H}_{\gamma}(\log R(x))=e^{-\gamma\log R(x)}\log R(x)$. As described earlier, the transform ${\mathcal{H}}_{\gamma}$ maps $\log R^{\prime}(s)={\mathcal{H}}_{\gamma}(\log R(s)),\forall s\in{\mathcal{S}}$ such that $\log R^{\prime}(s)$ is skewed towards higher reward states as a function of some parameter $\gamma$ (see §C.3 for further details. Note, ${\mathcal{H}}_{\gamma}$ is monotonic for values of $\log R(x)\leq 0$ and $\gamma\geq 0$. In our graph and sequence generations tasks, $\log R(x)\leq 0,\forall x\in{\mathcal{X}}$, hence we are able to use ${\mathcal{H}}_{\gamma}$ as a monotonic transform for $\log R(x)$. We show an example of ${\mathcal{H}}_{\gamma}$ applied to the cliques task in Figure 8 for various value sof $\gamma$.
@@ C.4 Evaluation metrics
To measure the generative modeling performance of GFlowNets we consider the Jensen-Shannon (JS) divergence and the MAE between the learned $\log p(x;\theta)$ and the true $\log p(x)$. For JS divergence, we compute:
@@
where $Q=\frac{1}{2}(p(x)+p(x;\theta))$. For the MAE, we simply take the absolute error $|\log p(x;\theta)-\log p(x)|$ and average over the cardinality of the state space.
@@ C.5 Model architectures
For all graph experiments we use a modified graph transformer (Veličković et al., 2017; Shi et al., 2020). On top of the normal attention mechanism, we augment the input of each layer with the output of one round of message passing (using layers from the work of Li et al., 2020) with a sum aggregation–we found that this was useful in tasks where counting was required. We use 8 layers with 128-dimensional embeddings and 4 attention heads; we use this architecture after having tried different numbers of layers and embeddings in order to validate our task; this is coincidentally represented in Fig. 1(a).
@@
We use the Adam optimizer with learning rate $0.0001$ for all models.